# Synthetic Object Compositions for Scalable and Accurate Learning in Detection, Segmentation, and Grounding

### v3

@@ 4.5 Task 5: Intra‐class referring expression, a challenging visual grounding task
Figure 6: (Section 4.5) Intra-class referring demands fine-grained attribute discrimination among same-category instances (e.g., picking out the "charcoal-grey cat"). Augmenting with SOS synthetic pairs (+SOC-FC-50K+SOC-GC-50K) yields the largest improvements in both metrics.
@@
Copy-Paste [Ghiasi2020SimpleCI] 9.32 35.2
X-Paste [Zhao2022XPasteRS] 9.41 37.2
SegGen [ye2024seggen] 9.73 36.8
SynGround [he2024learningsyntheticdatavisual] 40.1 / 89.2
@@
Intra‐class Referring Expression. Current visual grounding benchmarks—such as RefCOCO—can be circumvented. For example, to locate a “green car,” a model might ignore the attribute “green” and rely solely on the noun “car.” This shortcut breaks down when multiple cars of different colors and makes are present in the scene. We call this scenario intra-class referring, since it requires fine-grained attribute discrimination and is therefore more challenging. To systematically evaluate this use case, we curate an “intra-class referring expression” benchmark from COCO and OpenImages V7 [openimages] by selecting around 100 images that contain multiple instances of the same category exhibiting distinct attributes (colour, shape/pose, subtype). Each instance is tightly boxed and manually labeled with its key attribute variation (e.g. “red” vs. “blue”, “round” vs. “elongated”), yielding per‐instance category labels, bounding‐box coordinates, and fine‐grained attribute annotations. We propose two metrics for this new task: Average Gap: For each image, we compare the model’s confidence of the correct box (IoU $\geq 0.5$) against the highest-scoring same-category distractor. The gap between these two confidence scores measures how much more certain the model is about the true object than about any distractor. We report the average of this margin across all images; Positive Gap Ratio: We also calculate the percentage of images in which the ground-truth box receives the highest confidence score of all same-category candidates. This tells us in how many cases the model correctly prioritizes the intended instance over its distractors.
@@ Appendix B Detailed Related work
Open-Vocabulary detection and segmentation. Traditional object detection [girshick2015fast, ren2016faster] and segmentation [he2017mask, cheng2022maskedattentionmasktransformeruniversal] methods typically target predefined categories from standard benchmarks (e.g., COCO [lin2015microsoftcococommonobjects], OpenImages [Kuznetsova_2020], Object365 [9009553], LVIS [gupta2019lvisdatasetlargevocabulary]). Transformer-based architectures, such as DETR [Carion2020EndtoEndOD] and its derivatives [Zhu2020DeformableDD, Zhang2022DINODW, Lv2023DETRsBY, Huang2024DQDETRDW], have revolutionized end-to-end detection. Recent advances in vision-language models trained via contrastive language supervision [radford2021learning, Jia2021ScalingUV] have expanded detection to open-vocabulary settings [Zareian2020OpenVocabularyOD]. Approaches include distillation of pretrained vision-language knowledge into region-level representations [Gu2021OpenvocabularyOD, Zhong2021RegionCLIPRL], adapting contrastive models for detection [Minderer2022SimpleOO], integrating language-conditioned queries into DETR framework [Zang2022OpenVocabularyDW], and leveraging abundant image-level annotations to overcome detection annotation scarcity [Zhou2022DetectingTC, Lin2022LearningOA]. Building upon these approaches, scaling pre-training datasets has significantly advanced open-vocabulary models [li2021grounded, Yao2022DetCLIPDV, Yao2023DetCLIPv2SO]. GLIP [li2021grounded] combines detection with grounded language-image pretraining, introducing the ODInW benchmark [li2021grounded, li2022elevater]. Datasets like V3Det [wang2023v3detvastvocabularyvisual], with over 13,000 categories, further push open-vocabulary capabilities. MM-Grounding-DINO [Liu2023GroundingDM, zhao2024open], a combination of DINO [Zhang2022DINODW] and grounded pretraining, represents state-of-the-art performance alongside real-time solutions like Yolo-World [Cheng2024YOLOWorldRO]. Similarly, open-vocabulary segmentation (OVS) models [Zhou2021ExtractFD, ghiasi2022scalingopenvocabularyimagesegmentation, Li2022LanguagedrivenSS, Liang2022OpenVocabularySS] now incorporate language supervision, evaluated on standard semantic segmentation datasets (COCO-Stuff [Caesar2016COCOStuffTA], ADE20K [8100027, zhou2018semanticunderstandingscenesade20k], PASCAL VOC [everingham2011pascal], PASCAL Context [Mottaghi2014TheRO], and Cityscapes [Cordts2016Cityscapes]). Universal models like APE [shen2023aligningpromptinguniversalvisual] unify grounding, segmentation, and detection under diverse prompts [Zou2022GeneralizedDF, Zou2023SegmentEE].
@@
Referring expression and Grounding. Referring expression comprehension involves localizing image objects using natural language descriptions [Yu2016ModelingCI]. Early datasets, such as ReferItGame [kazemzadeh-etal-2014-referitgame], evolved into standard benchmarks like RefCOCO, RefCOCO+, and RefCOCOg [Yu2016ModelingCI, Mao2015GenerationAC], annotated over COCO images. Flickr30K Entities [plummer2016flickr30kentitiescollectingregiontophrase] and Visual Genome [Krishna2016VisualGC] facilitate visual grounding with detailed region annotations. Earlier methods employed modular decomposition [Hu2016ModelingRI] (e.g., subject and relationships detected individually in MattNet [Yu2018MAttNetMA]), but transformer-based models now dominate [Su2019VLBERTPO, li2019visualbertsimpleperformantbaseline, lu2019vilbertpretrainingtaskagnosticvisiolinguistic, chen2020uniteruniversalimagetextrepresentation, deng2022transvgendtoendvisualgrounding, Kamath2021MDETRM]. MDETR introduced a modulated detector conditioned directly on textual queries [Kamath2021MDETRM]. Recent open-vocabulary models, including GLIP and MM-Grounding-DINO, have also been adapted to this task [Huang2024DQDETRDW, li2021grounded, Liu2023GroundingDM, zhao2024open, deng2022transvgendtoendvisualgrounding, du2022visualgroundingtransformers]. Concurrently, the introduction of datasets like GoldG [li2022groundedlanguageimagepretraining] provides essential human annotations for grounding tasks. Multimodal large language models (MLLMs) have achieved state-of-the-art REC performance by directly generating bounding box coordinates from text [Wu2022GRiTAG, You2023FerretRA, Ma2024GromaLV]. Recent efforts like KOSMOS-2 [Peng2023Kosmos2GM] leverage synthetic data to overcome annotation scarcity, although verifying synthetic annotations remains challenging [Yao2024DetCLIPv3TV].
@@
Synthetic datasets for detection, segmentation, and visual grounding. Synthetic datasets have emerged to provide large-scale annotations without manual labeling, with early examples such as SYNTHIA [Ros2016TheSD] and GTA5 [Richter2016PlayingData] offering useful segmentation labels via game-engine rendering but revealed domain gaps when applied to real-world images. Virtual KITTI 2 [Cabon2020VKITTI2] and Synscapes [Wrenninge2018Synscapes] significantly improved realism for driving scenes through photorealistic and physically-based rendering, and Hypersim [Roberts_2021_ICCV] offered extensive indoor data with dense segmentation, depth, and normal annotations. More recently, diffusion-based synthesis has produced high-resolution, photorealistic images with controllable layouts [Rombach2021HighResolutionIS, Nichol2021GLIDETP, Ramesh2022HierarchicalTI, Li2023GLIGENOG], realistic illumination variations [Kocsis2024LightItIM, zhang2025scaling], and scalability to open-vocabulary and pseudo-labeled data [He2022IsSD, Azizi2023SyntheticDF, lee2025molmoactactionreasoningmodels, Fan2023ScalingLO, zheng2025brokentokenslanguagemodel, Zhang2023AddingCC, Jiang2022PseudoQGP, Peng2023Kosmos2GM, du2025paintoutsideboxsynthesizing, gao2024generatesceneevaluatingimproving, zhang2024provision].
@@
Early efforts at synthetic data relied heavily on 3D simulators, which require painstakingly crafted assets and still struggle to capture the diversity of real-world scenes. For purely 2D approaches, one branch resorts to simple copy-and-paste augmentations, transplanting segmented objects into real images [Dvornik2018ModelingVC, Fang2019InstaBoostBI, Ghiasi2020SimpleCI]. However, such methods remain constrained by the quality of available segments and produce conspicuous seams around pasted regions, causing models to exploit edge cues rather than true object semantics. X-Paste [Zhao2022XPasteRS] improves segment integrity and diversity by using diffusion models to enrich object segments before pasting, but it still randomly places segments and therefore does not fully resolve realism or seam-artifact issues. Moreover, because these methods are defined as augmentation pipelines applied to existing images, they are limited by the scale of the real data and restricted to a narrow set of categories (e.g., only 1.3 K categories on LVIS and COCO).
@@
A complementary line of work generates entire scenes with diffusion models or directly using real images, but applies off-the-shelf detectors to harvest pseudo-labels for bounding boxes. Wang et al. [Wang2024LearningVG] automatically produce referring annotations using VLMs. He et al. [he2024learningsyntheticdatavisual] leverage diffusion models to create varied images and use detectors to generate bounding-box labels. While these approaches achieve higher image fidelity, their pseudo-labels are noisy and yield smaller performance gains compared to real-image annotations, and also limit them to only high-level grounding tasks, instead of dense segmentation and detection tasks.
@@ D.1 Comparison with Synthetic Object Segments
Subjects 200K [tan2024omini] contains roughly 200K FLUX‑generated segments, but they cover <1K categories and are captured from only a handful of canonical viewpoints. Because the library offers neither broad category coverage nor multi‑view variation, it limits the semantic and geometric diversity attainable when composing new scenes.
@@ D.2 Comparison with Synthetic Data Pipelines
Simulator‑rendered datasets—SYNTHIA [Ros2016TheSD], GTA5 [Richter2016PlayingData], Virtual KITTI 2 [Cabon2020VKITTI2], Synscapes [Wrenninge2018Synscapes], and the indoor‑focused Hypersim [Roberts_2021_ICCV]—deliver pixel‑accurate masks and boxes by design, yet they remain confined to pre‑built domains (mostly driving or indoor scenes) and a closed object vocabulary; a noticeable photo‑to‑sim gap still emerges when models are applied to real photographs.
Copy‑paste families such as Simple Copy‑Paste [Ghiasi2020SimpleCI], InstaBoost [Fang2019InstaBoostBI], and the diffusion‑refined X‑Paste [Zhao2022XPasteRS] transplant real segments into new images. They are inexpensive and controllable, but the pasted objects often betray seam artefacts or lighting clashes. Because the pipeline augments a fixed real corpus, its scale is capped by the number of host images and its category list rarely exceeds ~1.3 K classes from LVIS/COCO.
Diffusion‑plus‑pseudo‑label pipelines invert the recipe. Methods like SynGround [he2024learningsyntheticdatavisual] and Learning VG [Wang2024LearningVG] first synthesise entire scenes with text‑to‑image models [Rombach2021HighResolutionIS, Nichol2021GLIDETP, Ramesh2022HierarchicalTI], then harvest boxes or phrases via detectors or VLMs. Although the images are photorealistic and open‑vocabulary, the labels inherit detector noise and therefore suit only coarse grounding rather than dense segmentation.
@@ D.3 Visual Realism: FID Score Analysis
To quantify the visual realism of our synthetic images, we computed Fréchet Inception Distance (FID) [Heusel2017GANsTB] scores on a 1K-sample subset comparing SOC against Copy-Paste and X-Paste baselines. As shown in Table 9, SOC achieves an FID of 131.93, substantially lower than both Simple Copy-Paste (165.55) and X-Paste (166.03). This indicates that even without explicitly optimizing for photorealism, SOC produces images that are more distributionally aligned with natural image statistics than alternative synthetic pipelines.
@@ E.1 Details of Mask-Area-Weighted Blending Algorithm
To achieve realistic lighting, we adopt IC‑Light [zhang2025scaling], a foreground‑conditioned diffusion model that ingests an image containing foreground objects, generates a matching background with a text prompt, and relights the composite to achieve photorealism.
@@ Appendix G Real Segments Filtering Pipeline
We annotate 4,000 samples—with GPT‑4o‐assisted chain‑of‑thought prompts—for ground truth and train three ViT‑B/16 classifiers, one per dimension [dosovitskiy2021imageworth16x16words]. At inference, we average the predicted scores and retain the top 30% of segments. We use this pipeline to filter the SA-1B, COCO, VOC, ADE20K, and get a total of 10M of real object segments for future exploration.

### v4

@@ 4.5 Task 5: Intra‐class referring expression, a challenging visual grounding task
Figure 6: (Section 4.5) Intra-class referring demands fine-grained attribute discrimination among same-category instances (e.g., picking out the "charcoal-grey cat"). Augmenting with SOC synthetic pairs (+SOC-FC-50K+SOC-GC-50K) yields the largest improvements in both metrics.
@@
Copy-Paste [20] 9.32 35.2
X-Paste [112] 9.41 37.2
SegGen [102] 9.73 36.8
SynGround [30] 40.1 / 89.2
@@
Intra‐class Referring Expression. Current visual grounding benchmarks—such as RefCOCO—can be circumvented. For example, to locate a “green car,” a model might ignore the attribute “green” and rely solely on the noun “car.” This shortcut breaks down when multiple cars of different colors and makes are present in the scene. We call this scenario intra-class referring, since it requires fine-grained attribute discrimination and is therefore more challenging. To systematically evaluate this use case, we curate an “intra-class referring expression” benchmark from COCO and OpenImages V7 [45] by selecting around 100 images that contain multiple instances of the same category exhibiting distinct attributes (colour, shape/pose, subtype). Each instance is tightly boxed and manually labeled with its key attribute variation (e.g. “red” vs. “blue”, “round” vs. “elongated”), yielding per‐instance category labels, bounding‐box coordinates, and fine‐grained attribute annotations. We propose two metrics for this new task: Average Gap: For each image, we compare the model’s confidence of the correct box (IoU $\geq 0.5$) against the highest-scoring same-category distractor. The gap between these two confidence scores measures how much more certain the model is about the true object than about any distractor. We report the average of this margin across all images; Positive Gap Ratio: We also calculate the percentage of images in which the ground-truth box receives the highest confidence score of all same-category candidates. This tells us in how many cases the model correctly prioritizes the intended instance over its distractors.
@@ Appendix B Detailed Related work
Open-Vocabulary detection and segmentation. Traditional object detection [22, 78] and segmentation [27, 8] methods typically target predefined categories from standard benchmarks (e.g., COCO [59], OpenImages [47], Object365 [84], LVIS [25]). Transformer-based architectures, such as DETR [5] and its derivatives [122, 108, 114, 34], have revolutionized end-to-end detection. Recent advances in vision-language models trained via contrastive language supervision [76, 36] have expanded detection to open-vocabulary settings [107]. Approaches include distillation of pretrained vision-language knowledge into region-level representations [23, 117], adapting contrastive models for detection [67], integrating language-conditioned queries into DETR framework [106], and leveraging abundant image-level annotations to overcome detection annotation scarcity [121, 58]. Building upon these approaches, scaling pre-training datasets has significantly advanced open-vocabulary models [56, 100, 99]. GLIP [56] combines detection with grounded language-image pretraining, introducing the ODInW benchmark [56, 51]. Datasets like V3Det [90], with over 13,000 categories, further push open-vocabulary capabilities. MM-Grounding-DINO [63, 113], a combination of DINO [108] and grounded pretraining, represents state-of-the-art performance alongside real-time solutions like Yolo-World [9]. Similarly, open-vocabulary segmentation (OVS) models [120, 21, 50, 57] now incorporate language supervision, evaluated on standard semantic segmentation datasets (COCO-Stuff [4], ADE20K [118, 119], PASCAL VOC [16], PASCAL Context [68], and Cityscapes [10]). Universal models like APE [85] unify grounding, segmentation, and detection under diverse prompts [124, 125].
@@
Referring expression and Grounding. Referring expression comprehension involves localizing image objects using natural language descriptions [105]. Early datasets, such as ReferItGame [42], evolved into standard benchmarks like RefCOCO, RefCOCO+, and RefCOCOg [105, 66], annotated over COCO images. Flickr30K Entities [71] and Visual Genome [46] facilitate visual grounding with detailed region annotations. Earlier methods employed modular decomposition [33] (e.g., subject and relationships detected individually in MattNet [104]), but transformer-based models now dominate [86, 53, 64, 6, 11, 40]. MDETR introduced a modulated detector conditioned directly on textual queries [40]. Recent open-vocabulary models, including GLIP and MM-Grounding-DINO, have also been adapted to this task [34, 56, 63, 113, 11, 13]. Concurrently, the introduction of datasets like GoldG [54] provides essential human annotations for grounding tasks. Multimodal large language models (MLLMs) have achieved state-of-the-art REC performance by directly generating bounding box coordinates from text [94, 103, 65]. Recent efforts like KOSMOS-2 [70] leverage synthetic data to overcome annotation scarcity, although verifying synthetic annotations remains challenging [101].
@@
Synthetic datasets for detection, segmentation, and visual grounding. Synthetic datasets have emerged to provide large-scale annotations without manual labeling, with early examples such as SYNTHIA [82] and GTA5 [79] offering useful segmentation labels via game-engine rendering but revealed domain gaps when applied to real-world images. Virtual KITTI 2 [3] and Synscapes [93] significantly improved realism for driving scenes through photorealistic and physically-based rendering, and Hypersim [80] offered extensive indoor data with dense segmentation, depth, and normal annotations. More recently, diffusion-based synthesis has produced high-resolution, photorealistic images with controllable layouts [81, 69, 77, 55], realistic illumination variations [44, 111], and scalability to open-vocabulary and pseudo-labeled data [29, 1, 49, 17, 115, 110, 37, 70, 14, 19, 109].
@@
Early efforts at synthetic data relied heavily on 3D simulators, which require painstakingly crafted assets and still struggle to capture the diversity of real-world scenes. For purely 2D approaches, one branch resorts to simple copy-and-paste augmentations, transplanting segmented objects into real images [15, 18, 20]. However, such methods remain constrained by the quality of available segments and produce conspicuous seams around pasted regions, causing models to exploit edge cues rather than true object semantics. X-Paste [112] improves segment integrity and diversity by using diffusion models to enrich object segments before pasting, but it still randomly places segments and therefore does not fully resolve realism or seam-artifact issues. Moreover, because these methods are defined as augmentation pipelines applied to existing images, they are limited by the scale of the real data and restricted to a narrow set of categories (e.g., only 1.3 K categories on LVIS and COCO).
@@
A complementary line of work generates entire scenes with diffusion models or directly using real images, but applies off-the-shelf detectors to harvest pseudo-labels for bounding boxes. Wang et al. [91] automatically produce referring annotations using VLMs. He et al. [30] leverage diffusion models to create varied images and use detectors to generate bounding-box labels. While these approaches achieve higher image fidelity, their pseudo-labels are noisy and yield smaller performance gains compared to real-image annotations, and also limit them to only high-level grounding tasks, instead of dense segmentation and detection tasks.
@@ D.1 Comparison with Synthetic Object Segments
Subjects 200K [88] contains roughly 200K FLUX‑generated segments, but they cover <1K categories and are captured from only a handful of canonical viewpoints. Because the library offers neither broad category coverage nor multi‑view variation, it limits the semantic and geometric diversity attainable when composing new scenes.
@@ D.2 Comparison with Synthetic Data Pipelines
Simulator‑rendered datasets—SYNTHIA [82], GTA5 [79], Virtual KITTI 2 [3], Synscapes [93], and the indoor‑focused Hypersim [80]—deliver pixel‑accurate masks and boxes by design, yet they remain confined to pre‑built domains (mostly driving or indoor scenes) and a closed object vocabulary; a noticeable photo‑to‑sim gap still emerges when models are applied to real photographs.
Copy‑paste families such as Simple Copy‑Paste [20], InstaBoost [18], and the diffusion‑refined X‑Paste [112] transplant real segments into new images. They are inexpensive and controllable, but the pasted objects often betray seam artefacts or lighting clashes. Because the pipeline augments a fixed real corpus, its scale is capped by the number of host images and its category list rarely exceeds ~1.3 K classes from LVIS/COCO.
Diffusion‑plus‑pseudo‑label pipelines invert the recipe. Methods like SynGround [30] and Learning VG [91] first synthesise entire scenes with text‑to‑image models [81, 69, 77], then harvest boxes or phrases via detectors or VLMs. Although the images are photorealistic and open‑vocabulary, the labels inherit detector noise and therefore suit only coarse grounding rather than dense segmentation.
@@ D.3 Visual Realism: FID Score Analysis
To quantify the visual realism of our synthetic images, we computed Fréchet Inception Distance (FID) [32] scores on a 1K-sample subset comparing SOC against Copy-Paste and X-Paste baselines. As shown in Table 9, SOC achieves an FID of 131.93, substantially lower than both Simple Copy-Paste (165.55) and X-Paste (166.03). This indicates that even without explicitly optimizing for photorealism, SOC produces images that are more distributionally aligned with natural image statistics than alternative synthetic pipelines.
@@ E.1 Details of Mask-Area-Weighted Blending Algorithm
To achieve realistic lighting, we adopt IC‑Light [111], a foreground‑conditioned diffusion model that ingests an image containing foreground objects, generates a matching background with a text prompt, and relights the composite to achieve photorealism.
@@ Appendix G Real Segments Filtering Pipeline
We annotate 4,000 samples—with GPT‑4o‐assisted chain‑of‑thought prompts—for ground truth and train three ViT‑B/16 classifiers, one per dimension [12]. At inference, we average the predicted scores and retain the top 30% of segments. We use this pipeline to filter the SA-1B, COCO, VOC, ADE20K, and get a total of 10M of real object segments for future exploration.